10. A 2.5 μF capacitor is charged to 250 V and then connected across a 1000 Ω resistor. The initial current just after the capacitor is connected to the resistor is a) 0 A c) 0.25 A b) 100 A d) infinity
Answer #1

Current just after connection of charged capacitor to resistor will be

i = V/R = 250/1000 = 0.25 A

At that moment the capacitor will behave just as a battery.

Answer is c) 0.25 A